identify pattern and next number
3,5,6,10,12,20

Answer by checkley77(12844)

3 5 6 10 12 20 the multipliers are 1 2/3 & 1.2
3(1 2/3)=5(1.2)=6(1 2/3)=10(1.2)=10(1 2/3)=20(1.2)=24 ans. for the next number.